OpenStudy (nannie):

For the problem x²-5x+10=0 Should I leave the answer as {5±√(25-40)}/2 or 2.5, 1.936i and 2.5, -1936i

OpenStudy (anonymous):

should have two answers + and negative

OpenStudy (anonymous):

x^2-5x+10 = 7.73 / -2.26

OpenStudy (jamesj):

If would leave them as \[\frac{5}{2} \pm \frac{\sqrt{15}}{2} i\] As a general rule, most mathematicians wouldn't convert this to a decimal because you lose accuracy by stopping at n decimal points, whatever n is; but the expression above is exact.